If 1+sin a=9 cos a 
then 5(1-cos a)=
  1. sin a
  2. 2 sin a
  3. 4 sin a 
  4. tan a

Grade:11

1 Answers

Vikas TU
7 years ago
1+sin a=9 cos a 
Divide by cosa both sides, we get
seca + tana = 9
seca = 9 – tana
squaring both sides,
(seca)^2 = 81 + (tana)^2  – 18tana
1 + (tana)^2  = 8` + tana^2 – 18tana
tan a = 7/18
cosa = 18/root(49 + 18^2)
sina = 7/root(49 + 18^2)
now cross check with all the options.